### v1

# Theoretical analysis of photon detection mechanism in superconducting single-photon detectors

Abstract
To elucidate the photon detection mechanism of superconducting single-photon detectors, we theoretically examine the dynamics of type-II superconductors with a bias current using the two-dimensional time-dependent Ginzburg-Landau and the Maxwell equations. The photon injection that weakens the superconducting order parameter is treated phenomenologically as a local temperature increase, and the amount of injection is controlled by the initial hotspot radius. The photon is detected by the voltage change between two electrodes attached to the left and right edges of the superconductor. We find that certain parameter ranges can be explained by the traditionally considered hotspot model, while other parameter ranges are governed by the generation and annihilation of superconducting vortex and antivortex pairs. The photon detection is possible for an initial hotspot radius that exceeds a threshold value. We find that the generation of a vortex–antivortex pair occurs near the threshold. The flow of the pair perpendicular to the current direction finally creates a normal region for the photon detection. The voltage change for the Ginzburg–Landau parameter close to the transition point from type-II to type-I superconductor shows anomalous behavior that is not associated with the dynamics of the vortex–antivortex pair. We also examine the effects of spatially non-uniform current density on the voltage change and the superconducting order parameter to provide a hint to understand the behavior of wide-strip single-photon detectors. The estimated values of incident photon energy and response time for photon detection are reasonable in comparison with experiments. The present comprehensive examination provides useful guidelines for flexible design of device structures.
I Introduction
The superconducting single-photon detector (SSPD) has many advantages, such as fast response, detection efficiency, low jitter, and low dark count [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. The extremely high photon detection capability is essential for photonic quantum information technologies. Nevertheless, the photon detection mechanism of the SSPD is not adequately understood. Therefore, currently there are no specific guidelines for the selection of superconducting materials and the design of devices to further improve performance. Typically, SSPD devices are composed of a superconducting nanostrip with a meander pattern (superconducting nanowire single-photon detector, SNSPD), and there are still restrictions regarding the polarization dependence and productivity. In this paper, our aim is to investigate the SSPD mechanism to overcome this situation.
The SNSPD mechanism has been widely believed to be described by a hotspot model. According to this model, the device size for high-performance detection is severely limited, and for this reason the nanostrip structure is usually fabricated. However, recent experiments have shown that single-photon detection can also be performed with a micrometer-sized device [12, 13] (superconducting wide-strip photon detector, SWSPD). In particular, the authors in Ref. [13] propose a high critical current bank structure to suppress the intrinsic dark count caused by the nonuniform distribution of superconducting currents in the microstrip. The experimental successes for efficient photon detection by using the wide strip provide a good opportunity to reconsider the traditional hotspot model for the SSPD and SNSPD mechanisms.
In the hotspot model, the superconductivity is locally suppressed by the incident photon whose energy scale is two or three orders of magnitude larger than the superconducting gap. In this case, without going into the details of the microscopic electronic states, it can be expected that the suppression is a thermal process. We refer to this process as initial hotspot formation. The superconducting current in the SSPD device flows outside the hotspot region. If the device is composed of a thin wire, the increase in the superconducting current density outside the hotspot region may exceed a critical value, leading to local breaking of the superconducting state. This process can be observed by the voltage between the two wire terminals. Although the formation of the initial hotspot region is plausible, the description of the subsequent processes seems to be insufficient to understand the detection mechanism in the micrometer-sized SSPD devices. In the subsequent processes, the hotspot model focuses on the expansion of the region, and there are many previous studies on the hotspot dynamics in terms of the heat-diffusion equation [14, 15, 16, 17]. However, analyzing the thermal distribution alone is not sufficient for a detailed understanding of electromagnetic response.
An alternative possible mechanism is based on the formation of the vortex–antivortex pair. This idea was originally proposed in Refs. [18, 19], but to the best of our knowledge, no detailed study has been reported so far. In this vortex–antivortex model, the dynamics of vortices and antivortices perpendicular to the current flow direction result in a voltage change. A vortex-assisted photon-count mechanism was also proposed in Ref. [20], inspired by the single-vortex-crossing mechanism for dark counts [21]. A recent study [22] reports a unified theory of dark‐count rate and system‐detection efficiency based on a vortex-crossing framework. The photon detection mechanism based on this model is expected to be less sensitive to the finite-size limitations, consistent with the realization of SWSPD. The creation of vortex–antivortex pairs is also observed by the local heating effect of a scanning tunneling microscope tip [23]. Theoretical confirmation of pair generation would represent an important step toward developing a comprehensive understanding of the physics of SSPD devices. At the same time, we aim to clarify how the model parameters differ between the hotspot and vortex–antivortex models.
Here, we summarize two previous theoretical works that are based on the time-dependent Ginzburg–Landau (TDGL) equation and are closely related to the present work.
[…]
The examination of the vortex–antivortex mechanism in terms of the TDGL equation is reported in Ref. [25]. The authors numerically solved the two-dimensional TDGL equation coupled with the Poisson equation for the electric potential and the heat-diffusion equation. They do not take into account the vector potential for the magnetic field. This assumption is valid in the large-$\kappa$ limit. They take the coherence length at zero temperature, the initial radius of the hotspot, and the width of the film (perpendicular to the current-flow direction) as $\xi_{0}=7.5$ nm, $R_{\mathrm{init}}=9$ nm, and $w=13\xi_{0}\sim 52\xi_{0}$, respectively. The local temperature increase is determined by energy conservation $E_{\mathrm{photon}}=\pi R_{\mathrm{init}}^{2}dC_{v}\Delta T$, where $C_{v}$ is a heat capacity, $d$ is a thickness of the sample, and $\Delta T$ is a temperature increase due to photon injection. At current larger than a threshold (the authors call this detecting current), they found that the superconducting state collapses starting from the appearance of a vortex–antivortex pair in the center of the initial hotspot region. Lorentz force causes their motion that heats the system locally and gives rise to a normal domain. The authors particularly focus on the dependence of detecting current on $\Delta T$ and $w$. In this study, $\xi_{0}$ is comparable to $R_{\mathrm{init}}$, but we consider that the initial hotspot size strongly affects the stability of the formation of a vortex–antivortex pair. Therefore, the parameters should be determined more carefully. Furthermore, the authors do not comment on why the traditional hotspot mechanism is not consistent with their analysis.
Motivated by the present status mentioned above, we examine the dynamics of type-II superconductors with a bias current using the two-dimensional TDGL and the Maxwell equations. The photon injection is treated phenomenologically by the temperature increase inside the initial hotspot. The important parameter in this study is the initial hotspot radius by which the amount of photon energy is controlled. We find that certain parameter ranges can be explained by the hotspot model, while other parameter ranges are governed by the generation and annihilation of a vortex–antivortex pair. The photon detection is possible for an initial hotspot radius that exceeds a threshold value. We demonstrate that the pair generation occurs near the threshold. The flow of the pair perpendicular to the current direction finally leads to the formation of a normal region responsible for photon detection. During the formation process, the voltage change shows oscillation that originates from repeated penetration of vortices and antivortices from the sample edges. The voltage change for the GL parameter close to the transition point from type-II to type-I superconductor shows anomalous oscillations that are not associated with the vortex–antivortex pair. We also examine the effects of spatially non-uniform current density on the voltage change and the superconducting order parameter in order to understand the behavior of wide-strip single-photon detectors. The estimated values of incident photon energy and response time for photon detection are reasonable in comparison with experiments. The present comprehensive examination provides useful guidelines for flexible design of device structures.
The organization of this paper is as follows. In Sect. II, we introduce our model and the calculation method. In Sect. III, we show our numerical results. Finally, we summarize our conclusion in Sect. IV.
II Model
Figure 1: Illustration of our system. The red and blue areas correspond to superconductor and metallic electrodes, respectively. The system size and current-flow direction are indicated, respectively. The superconducting region is given by $0\leq x\leq L_{x}$ and $0\leq y\leq L_{y}$. The origin of the initial hotspot with radius $R_{\mathrm{init}}$ is set to the center of the superconductor.
Figure 1 illustrates our system. The system is composed of a two-dimensional superconductor at zero temperature ($T_{\mathrm{sys}}=0$) with a size of $L_{x}\times L_{y}=40\xi_{0}\times 20\xi_{0}$, where $\xi_{0}$ is the coherent length of the superconductor at zero temperature. In ultrathin film NbN with the film size of $4\sim 5$ nm, $\xi_{0}$ is about $5$ nm, and thus the width of the present system is about $100$ nm. The left and right terminals are connected to metallic electrodes and the upper and lower edges are connected to vacuum. The bias current $j_{\mathrm{bias}}$ is injected from the left electrode. In most of the numerical simulations, we take $\kappa=10\gg 1/\sqrt{2}$, and thus the system is in the type-II superconductor.
In order to study the dynamics of the order parameter and electromagnetic fields, we solve the TDGL and the Maxwell equations, which are given, respectively, by
$-\frac{\hbar^{2}}{2m^{\ast}D}\left(\frac{\partial}{\partial t}-i\frac{e^{\ast}}{\hbar}\varphi\right)\Delta$
$\;\;=-\alpha\left(T\left({\bf\it r}\right)\right)\Delta+\beta\left|\Delta\right|^{2}\Delta+\frac{1}{2m^{\ast}}\left(-i\hbar{\bf\it\nabla}+\frac{e^{\ast}}{c}{\bf\it A}\right)^{2}\Delta,$ (1)
${\bf\it\nabla}\times\left({\bf\it\nabla}\times{\bf\it A}\right)=\frac{4\pi}{c}{\bf\it j},$ (2)
with the current density $\bf\it j$} defined by
[…]
where the first and second terms of $\bf\it j$} correspond to superconducting and normal currents, denoted by ${\bf\it j}_{s}$ and ${\bf\it j}_{n}=\sigma{\bf\it E}$, respectively. Here, $\Delta$, $\varphi$, $\bf\it A$}, $D$, $\sigma$, $e^{\ast}(>0)$, and $m^{\ast}$ are the superconducting order parameter, scalar potential, vector potential, diffusion constant, electric conductivity in the normal region, and effective mass and effective charge of a superconducting electron, respectively. The parameter $\alpha$ is determined phenomenologically by $\alpha\left(T\left({\bf\it r}\right)\right)=\alpha(0)\left(1-T({\bf\it r})/T_{c}\right)$, where $T({\bf\it r})$ is the space-dependent temperature due to the formation of the initial hotspot region and subsequent dynamics, $T_{c}$ is the critical temperature and $\alpha_{0}$ is a positive constant. The parameter $\beta$ is a positive constant. By a phase redefinition of the order parameter and the gauge transformation of the vector potential $\bf\it A$}, we will eliminate the scaler potential $\varphi$.
We summarize the boundary conditions employed in our numerical simulations. The upper and lower edges of the superconducting sample are connected to vacuum, and the left and right edges of the superconducting sample are connected to metallic electrodes. The Neumann boundary condition is imposed at superconductor/vacuum interfaces:
${\bf\it e}_{y}\cdot\left.\left(-i\hbar{\bf\it\nabla}+\frac{e^{\ast}}{c}{\bf\it A}\right)\Delta\right|_{y=0,L_{y}}=0,$ (4)
where ${\bf\it e}_{\mu=x,y,z}$ is a unit vector along the $\mu$ direction. At the left and right edges of the superconductor, we consider the following condition for the order parameter:
$\left.\Delta\right|_{x=0,L_{x}}=0.$ (5)
We assume that the magnetic field, induced by the transport current, is along the $z$ direction, $\bm{B}=B_{z}\bm{e}_{z}=\bm{\nabla}\times\bm{A}$, and consider the following condition for $B_{z}$. At the upper (++) and lower (-) superconductor/vacuum interfaces, the magnetic fields are given by
$\left.B_{z}\right|_{y=0,L_{y}}=\pm\frac{L_{y}}{2}\frac{4\pi}{c}j_{\mathrm{bias}},$ (6)
where the bias current $j_{\mathrm{bias}}$ satisfies the following conditions:
$j_{\mathrm{bias}}$ $=\frac{1}{L_{y}}\int_{0}^{L_{y}}j_{x}(x=0,y)dy$
$=\frac{1}{L_{y}}\int_{0}^{L_{y}}\left(j_{n}\right)_{x}(x=0,y)dy.$ (7)
We assume a spatially uniform normal current at both the metallic electrodes [i.e., $j(x=0,y)=j(x=L_{x},y)=j_{\mathrm{bias}}$], and the boundary condition on the magnetic fields there is given by
$\left.B_{z}\right|_{x=0,L_{x}}=\frac{4\pi}{c}j_{\mathrm{bias}}y.$ (8)
In this study, the photon injection is treated phenomenologically. A simple method is to consider heating in a fixed area due to the photon injection. We introduce the following temperature difference in and out of the initial hotspot:
$T({\bf\it r})=\begin{cases}T_{\mathrm{sys}}+\Delta T&0\leq|{\bf\it r}-{\bf\it r}_{c}|\leq R_{\mathrm{init}}\\ T_{\mathrm{sys}}&|{\bf\it r}-{\bf\it r}_{c}|>R_{\mathrm{init}}\end{cases},$ (9)
where $R_{\mathrm{init}}$ is the radius of the circular hotspot initially created by photon injection, ${\bf\it r}_{c}=(L_{x}/2,L_{y}/2)$ denotes the center of the initial hotspot, and $\Delta T$ is the temperature increase by photon injection. We take $T_{\mathrm{sys}}=0$ and $\Delta T=T_{c}$.
Let us define the voltage change $\Delta V$ between the left and right terminals of the superconducting sample by the path integral of the longitudinal component of the electric field as $\Delta V=-\int_{\bm{r}_{R}}^{\bm{r}_{L}}\bm{E}_{L}\cdot d\bm{r}$, where $\bm{r}_{L}=\left(3\xi_{0},L_{y}/2\right)$ and $\bm{r}_{R}=\left(L_{x}-3\xi_{0},L_{y}/2\right)$. The measurement positions of $\Delta V$ are taken at distance $3\xi_{0}$ from the terminals to avoid the effects of contact resistance on $\Delta V$. Because $\bm{E}_{L}$ can be written with a scalar field $V$ as $\bm{E}_{L}=-\bm{\nabla}V$, the integral is independent of the path connecting $\bm{r}_{L}$ and $\bm{r}_{R}$, i.e., $\Delta V=V(\bm{r}_{L})-V(\bm{r}_{R})$. The equation for $V$ is given by
$\nabla^{2}V=-\bm{\nabla}\cdot{\bm{E}},$ (10)
with the boundary conditions
$\left.\dfrac{\partial V}{\partial x}\right|_{x=0,L_{x}}=-\dfrac{j_{\mathrm{bias}}}{\sigma},\quad\left.\dfrac{\partial V}{\partial y}\right|_{y=0,L_{y}}=0.$ (11)
Let us introduce the following normalization for convenience of numerical simulation: $\tilde{{\bf\it\nabla}}={\bf\it\nabla}/\xi_{0}$, $\tilde{t}=t/t_{0}$, $\tilde{T}=T/T_{c}$, ${\bf\it\tilde{A}}={\bf\it A}/\xi_{0}H_{c2}$, and $\tilde{\Delta}=\Delta/\Delta_{\infty}(0)$, where $t_{0}=\xi_{0}^{2}/(12D)=4\pi\sigma\lambda_{0}^{2}/c^{2}$, $H_{c2}=c\hbar/e^{\ast}\xi_{0}^{2}$ is the upper critical field, and $\Delta_{\infty}(0)=\sqrt{m^{\ast}c^{2}/(4\pi e^{\ast 2}\lambda_{0}^{2})}$ is the zero-temperature order parameter in the absence of the magnetic field. Note that using the parameters in the TDGL equation, $\Delta_{\infty}(0)=\sqrt{\alpha(0)/\beta}$, $\xi_{0}=\hbar/\sqrt{2m^{\ast}\alpha(0)}$, and $\lambda_{0}=\sqrt{m^{\ast}c^{2}/(4\pi e^{\ast 2})}(\beta/\alpha(0))$. A typical value of $t_{0}=\xi_{0}^{2}/12D$ for NbN is about $4.2\times 10^{-2}$ ps (we assume $\xi_{0}=5.0$ nm and $D=0.5$ cm2/s). Then, the TDGL and the Maxwell equations are normalized as follows:
$\partialderivative{\tilde{\Delta}}{\tilde{t}}=-\frac{1}{12}\left[\left(-i{\bf\it\tilde{\nabla}}+{\bf\it\tilde{A}}\right)^{2}\tilde{\Delta}-\left(1-\tilde{T}({\bf\it\tilde{r}})\right)\tilde{\Delta}+|\tilde{\Delta}|^{2}\tilde{\Delta}\right$], (12)
[…]
To preserve the gauge invariance of these equations even after discretization, we employ the link-variable method [26]. In the numerical simulation, the system is discretized into $N_{x}\times N_{y}$ meshes with a mesh spacing of $\delta=0.25\xi_{0}$ in both directions, i.e., $\delta=L_{x}/N_{x}=L_{y}/N_{y}$. For even integers $N_{x}$ and $N_{y}$, the numerical grid points are defined as $\bm{r}_{i_{x},i_{y}}=(i_{x}-1/2,i_{y}-1/2)\delta$, where $i_{\mu=x,y}=1,\cdots,N_{\mu}$. Note that there are no grid points located at $x=L_{x}/2$ or $y=L_{y}/2$ due to the above discretization; in the following calculations, these positions are given by $(L_{x}-\delta)/2$ and $(L_{y}-\delta)/2$, respectively; nevertheless, for simplicity, we denote them as $L_{x}/2$ and $L_{y}/2$, as long as no confusion arises. In particular, as shown below, such a shift in $\bm{r}_{c}$ leads to asymmetric dynamics. However, this asymmetry is not essential for photon detection, and we have confirmed that the symmetric dynamics are recovered when the hotspot center is set exactly at the system center, $\bm{r}_{c}=(L_{x}/2,L_{y}/2)$.
III Results
III.1 Dependence of $R_{\mathrm{init}}$ on voltage change
Let us present numerical results. We first examine the dependence of $R_{\mathrm{init}}$ on the photon detection performance for $j_{\mathrm{bias}}/j_{0}=0.28$ and $\kappa=10$. The bias current density is set to be slightly lower than the critical current density (the critical value of $j_{\mathrm{bias}}/j_{0}$ is about $0.29$), where $j_{0}=(e^{\ast}/m^{\ast})\Delta_{\infty}^{2}(0)(\hbar/\xi_{0})$. The photon detection performance is characterized by the time evolution of the voltage change $\Delta V$, which is normalized by $V_{0}=H_{c2}\xi_{0}^{2}/t_{0}$. Here, we do not consider any thermal relaxation processes in which $\Delta V$ again goes to zero toward the next photon detection, since we particularly focus on the detection mechanism. We believe that our conclusions do not change so much even if we additionally introduce these thermal relaxation processes.
Figure 2: Time evolution of voltage change $\Delta V$. Various lines present the profiles with different values of $R_{\mathrm{init}}/\xi_{0}$. We take $j_{\mathrm{bias}}/j_{0}=0.28$ and $\kappa=10$.
As shown in Fig. 2, for $R_{\mathrm{init}}/\xi_{0}\geq 2.9$, $\Delta V$ exhibits a rapid increase at a certain time and subsequently saturates, indicating successful photon detection. We refer to this time as the response time. For example, the response time is about $t/t_{0}=250$ for $R_{\mathrm{init}}/\xi_{0}=3.0$ (see green line in Fig. 2). The time for NbN is estimated as $t\sim 10$ ps. The response time depends on $R_{\mathrm{init}}$: a smaller value of $R_{\mathrm{init}}$ leads to a slower response time. For $R_{\mathrm{init}}/\xi_{0}>3.3$, the response time is smaller than $t/t_{0}=100$. On the other hand, for $R_{\mathrm{init}}/\xi_{0}<3.3$, the response starts to become noticeably slow. The magnitude of $\Delta V$ after saturation becomes smaller as $R_{\mathrm{init}}$ decreases. In Fig. 2, we also observe the oscillatory behavior of $\Delta V$ after the response time. This oscillation becomes more visible for smaller values of $R_{\mathrm{init}}$. We will address the origin of this oscillation in the next subsections.
Figure 3: Voltage change $\Delta V$ at $t/t_{0}=400$ (open squares) and $t/t_{0}=800$ (filled squares) as a function of the initial hotspot radius $R_{\mathrm{init}}$. We take $j_{\mathrm{bias}}/j_{0}=0.28$ and $\kappa=10$.
The trends in response time and the magnitude of $\Delta V$ as a function of $R_{\mathrm{init}}$ shown in Fig. 2 suggest that there exists a threshold value of $R_{\mathrm{init}}$ for photon detection. To see the threshold, we plot $\Delta V$ as a function of $R_{\mathrm{init}}$ in Fig. 3. All data are taken at $t/t_{0}=400$ and $t/t_{0}=800$. A sharp threshold is found: $\Delta V$ becomes finite for the parameter range $R_{\mathrm{init}}/\xi_{0}\geq 2.9$, whereas we have confirmed that, for $R_{\mathrm{init}}/\xi_{0}=2.8$, the resistive state is not attained even up to $t/t_{0}=1600$.
III.2 Dynamics of amplitude and phase of superconducting order parameter
Figure 4: Schematic view of a vortex–antivortex pair generated inside of the hotspot region as well as a penetrating vortex and a penetrating antivortex from outside the superconductor. Cross symbols inside circles indicate the current flow direction.
The oscillation of $\Delta V$ observed in Fig. 2 originates from the motion of a vortex–antivortex pair generated in the initial hotspot region and from their interactions with vortices and antivortices penetrating the superconductor from outside. Here, the latter votices and antivortices are called penetrating votices and penetrating antivortices, respectively, and they are distinguished from the vortex–antivortex pair generated in the hotspot region. Figure 4 is a schematic viewgraph of these vortices and antivortices.
Figure 5: Time evolution of spacetime profile of $\left|\tilde{\Delta}\left(x=L_{x}/2,y\right)\right|$ for $150\leq t/t_{0}\leq 500$. The blight region corresponds to stable superconducting state. The position, $y=10$, corresponds to the hotspot center. (a) $R_{\mathrm{init}}/\xi_{0}=3.0$, $j_{\mathrm{bias}}/j_{0}=0.28$, and $\kappa=10$. (b) $R_{\mathrm{init}}/\xi_{0}=3.8$, $j_{\mathrm{bias}}/j_{0}=0.25$, and $\kappa=5$.
Let us examine the origin of the oscillation in terms of the order parameter. For this purpose, we first present the time evolution of $\left|\tilde{\Delta}\left(x=L_{x}/2,y\right)\right|$ (passing through the hotspot area) in Fig. 5(a), where we take $R_{\mathrm{init}}/\xi_{0}=3.0$, $j_{\mathrm{bias}}/j_{0}=2.8$, and $\kappa=10$. For $t/t_{0}\leq 250$, there are no significant changes outside the initial hotspot region, but the dark region begins to develop inside the hotspot. This dark region flows along the $y$ axis and arrives at the upper and lower edges of the superconductor around $t/t_{0}=250$. After these initial dynamics, the dark and blight areas appear repeatedly in the intermediate region between the sample edges and the initial hotspot region. The repetition period is equal to the oscillation period observed in Fig. 2.
For comparison with the case where $\kappa$ is smaller, we present $\left|\tilde{\Delta}\left(x=L_{x}/2,y\right)\right|$ for $j_{\mathrm{bias}}/j_{0}=0.25$, $\kappa=5$, and $R_{\mathrm{init}}/\xi_{0}=3.8$ in Fig. 5(b). The oscillatory behavior of the dark and blight areas is also observed. However, before the flow of the dark area in the initial hotspot toward the edges, another dark region begins to penetrate from the lower edge of the superconductor around $t/t_{0}=350$. After that, these dark areas merge with each other. The dynamics of the (anti)vortex penetration is spatially asymmetric about $y=L_{y}/2$. This is due to the slight shift of the initial hotspot from the center of the system in the lattice discretization as mentioned at the end of Sect. II. We have confirmed numerically that this asymmetry does not occur when the hotspot is exactly located at the center of the system. As shown in the next paragraph, the dark regions originates in the vortex–antivortex pair generated in the initial hotspot for $t/t_{0}\lesssim 250$ and the penetrating vortices and antivortices for $t/t_{0}\gtrsim 250$.
Figure 6: Dynamics of amplitude [(a) left panels] and phase [(b) right panels] of superconducting order parameter. We take $t/t_{0}=155,222,238,250,324$ from top to bottom, $R_{\mathrm{init}}/\xi_{0}=3.0$, $j_{\mathrm{bias}}/j_{0}=0.28$, and $\kappa=10$. In the panels, contours are drawn every 0.1 for $3\xi_{0}\leq x\leq L_{x}-3\xi_{0}$; solid (dashed) lines indicate levels of 0.1, 0.3, 0.5, 0.7, and 0.9 (0.2, 0.4, 0.6, and 0.8).
To visualize the generation of the vortex–antivortex pair in the initial hotspot region and the formation of the normal region, we present color maps of the amplitude of the superconducting order parameter $\tilde{\Delta}(x,y)$ for various times in Fig. 6(a). We take $R_{\mathrm{init}}/\xi_{0}=3.0$, $j_{\mathrm{bias}}/j_{0}=0.28$, and $\kappa=10$. We find that the normal region is finally formed across the upper and lower ends of the sample, and this is consistent with the voltage change in Fig. 3. During this process, we observe two small dark spots generated in the initial hotspot and moving separately toward the ends of the sample [see panels (ii)–(iv) of Fig. 6(a)]. These spots are smaller than $R_{\mathrm{init}}/\xi_{0}=3$, and are comparable in size to the vortex core (of order $\xi_{0}$). After the dark spots reach the sample edges around $t/t_{0}=250$, the solid normal region is finally formed across the upper and lower edges [see Fig. 6(a)(v)]. Our observation cannot be simply understood in terms of the hotspot model.
We also plot the phase of the order parameter in Fig. 6(b). We clearly observe that the two small dark spots correspond to quantum vortices with opposite phase windings ($\pm 2\pi$ phase rotations) [see panels (ii)–(iv) of Fig. 6(b)]. Thus, we conclude that they are a vortex–antivortex pair. We find that the intermediate region between the two singularities expands as $t$ increases. At $t/t_{0}=250$ [see Fig. 6(b)(iv)], this region is almost in contact with the lower edge of the sample. According to Fig. 2, a sharp increase in $\Delta V$ occurs around this time. We cannot clearly observe the presence of penetrating vortices and antivortices at least up to $t/t_{0}\leq 250$. After $t/t_{0}=250$, $\Delta V$ starts oscillating, as shown in Fig. 2.
By careful analysis of the time evolution of the order parameter, we have confirmed repeated penetration of vortex-antivortex pairs from the outside of the superconductor, and the correspondence between the period of the oscillation in $\Delta V$ and that of the repeated penetration. For instance, the first local minimum and the second local maximum in $\Delta V$ are located around $t/t_{0}=254.5$ and $t/t_{0}=261$, respectively, and the vortex and antivortex penetrate from the outside around $t/t_{0}=258$. The time when $\Delta V$ reaches its local minimum corresponds to the onset of the decrease in the edge order parameter, serving as a precursor to the penetration of a vortex or an antivortex, while its local maximum is reached after the penetrating vortex and antivortex annihilate near the center.
III.3 Parameter region where the hotspot model is plausible
As mentioned in the previous subsections, the numerical results for smaller $R_{\mathrm{init}}$ values ($R_{\mathrm{init}}/\xi_{0}\sim 3.0$) and for $\kappa=10$ suggest the vortex–antivortex mechanism for photon detection. In Fig. 2, we have already found that the voltage change shows a rather smooth curve for larger $R_{\mathrm{init}}$. The absence of the oscillation may lead to a photon detection mechanism different from the vortex–antivortex pair dynamics. By focusing on the case of larger $R_{\mathrm{init}}$, we demonstrate a hotspot-like behavior for photon detection.
Figure 7: Dynamics of amplitude [(a) left panels] and phase [(b) right panels] of superconducting order parameter for $R_{\mathrm{init}}/\xi_{0}=6.0$, $j_{\mathrm{bias}}/j_{0}=0.28$, and $\kappa=10$. We take $t/t_{0}=35,38,55$ from top to bottom. In the left panels, the contour lines are plotted in the same manner as in Fig. 6(a).
Figure 7(a) presents the dynamics of the amplitude of $\tilde{\Delta}(x,y)$ for $R_{\mathrm{init}}/\xi_{0}=6.0$, $j_{\mathrm{bias}}/j_{0}=0.28$, and $\kappa=10$. In this case, $R_{\mathrm{init}}$ is comparable to the penetration depth ($\lambda_{0}/\xi_{0}=10$). We find that the hotspot region simply expands and touches the upper and lower edges of the superconductor at the initial stage of the time evolution [see panel (i) of Fig. 7(a) for $t/t_{0}=35$]. According to Fig. 2, $\Delta V$ already takes nearly half of its maximum value at $t/t_{0}=35$. Small dark spots with size $\xi_{0}$ cannot be observed. When $R_{\mathrm{init}}$ is large, the current does not circulate around the hotspot. Then, the vortex–antivortex pairs are less likely to form.
The phase of $\tilde{\Delta}(x,y)$ is also shown in Fig. 7(b). No $2\pi$ rotation of the phase occurs in the initial stage of the time evolution, although the normal region has almost been formed at $t/t_{0}=35$. At $t/t_{0}=38$ and $55$, we observe singularities in the phase. These should be interpreted with caution, since they are likely numerical artifacts arising from the fact that the order-parameter amplitude in the dark region has largely diminished, making it difficult to determine the phase accurately.
III.4 Voltage change and dynamics of order parameter near the transition point between type-I and type-II superconducting states
According to Ref. [24], where the authors assumed $\kappa=2$, they concluded that the simple hotspot expansion is a possible mechanism of photon detection. To compare it with the present work, we examine the dependence of $\kappa$ on $\Delta V$ and dynamics of the superconducting order parameter. We expect that the formation of the vortex–antivortex pairs becomes unstable as $\kappa$ decreases. As already shown in Fig. 4(b) for $\kappa=5$, we still observe the vortex–antivortex pair and the penetrating vortices and antivortices that have been observed for $\kappa=10$. We examine the case of $\kappa=1$, which is close to the transition ($\kappa_{c}=1/\sqrt{2}$) from type-II to type-I superconductor.
Figure 8: Time evolution of voltage change $\Delta V$. Various lines present the profiles with different values of $R_{\mathrm{init}}/\xi_{0}$. We take $j_{\mathrm{bias}}/j_{0}=0.09$ and $\kappa=1$.
Figure 8(a) shows the time evolution of $\Delta V$ for $\kappa=1$. We take $j_{\mathrm{bias}}/j_{0}=0.09$, which is slightly lower than the critical current. We observe the oscillatory behavior on $\Delta V$ especially for small $R_{\mathrm{init}}$. However, the profile is quite different from the previous one in Fig. 2 for $\kappa=10$: faster and slower oscillations coexist.
Figure 9: Dynamics of amplitude [(a) left panels] and phase [(b) right panels] of superconducting order parameter for $R_{\mathrm{init}}/\xi_{0}=5.6$, $j_{\mathrm{bias}}/j_{0}=0.09$, and $\kappa=1$. We take $t/t_{0}=330,430,500,550,800$ from top to bottom. In the left panels, the contour lines are plotted in the same manner as in Fig. 6(a).
To examine the difference, we show the dynamics of the order parameter for $R_{\mathrm{init}}/\xi_{0}=5.6$ in Fig. 9 (see red line in Fig. 8). In this case, $\Delta V$ shows a step-like structure with oscillation around $400\leq t/t_{0}\leq 550$. We find the presence of penetrating antivortices from the lower edge of the sample in the initial stage of time evolution [see Figs. 9(a)(ii) and 9(b)(ii) for $t/t_{0}=430$]. We also find the accumulation of penetrating antivortices in the initial hotspot area [see Figs. 9(a)(iii) and 9(b)(iii) for $t/t_{0}=500$]. The accumulation process occurs, when $\Delta V$ shows the step-like structure. The vortex–antivortex pair is not formed inside the hotspot. The oscillation of $\Delta V$ around $400\leq t/t_{0}\leq 550$ originates from penetration of antivortices at the lower edge. We have numerically confirmed that each penetration process occurs during the interval between the local minimum and local maximum of $\Delta V$. The next accumulation process starts by the penetration of vortices from the upper edge of the sample. The collision among these penetrated vortices and accumulated antivortices in the hotspot spikes the voltage.
The oscillation of $\Delta V$ is observed for both of $\kappa=10$ and $\kappa=1$, but the origins of the oscillation differ between these two cases.
III.5 Effects of spatially non-uniform current density on voltage change and dynamics of order parameter: a perspective for SWSPD
One of the major difficulties in achieving photon detection using wide superconducting strips is the presence of intrinsic dark count as a result of the non-uniform distribution of the superconducting current in the strip. To examine the effect of the non-uniformity on the photon detection performance, we would like to ask how our result changes when the current flow becomes spatially more non-uniform. For this purpose, we examine the influence of $L_{y}$, the sample width, on the dynamics of the superconducting order parameter.
Figure 10: Current density $j_{x}\left(x=L_{x}/2,y\right)$ before photon injection (dashed lines) and after the formation of the initial hotspot (solid lines) for $L_{x}\times L_{y}=40\xi_{0}\times 20\xi_{0}$ (red lines) and $L_{x}\times L_{y}=40\xi_{0}\times 40\xi_{0}$ (blue lines). Note that the integrated value of $j_{x}$ by $y$, $j_{\mathrm{bias}}$, is conserved. The current density after the hotspot formation is calculated at $t/t_{0}=150$. The initial hotspot radius is taken as $R_{\mathrm{init}}/\xi_{0}=3.0$ ($4.7$) for $L_{y}=20\xi_{0}$ ($40\xi_{0}$).
Figure 10 presents $j_{x}$ as a function of $y$ before and after photon injection. We consider a system with $L_{x}\times L_{y}=40\xi_{0}\times 40\xi_{0}$ which is twice the width of the other calculations (in this case we take $N_{y}=160$ with keeping linear mesh size $0.25\xi_{0}$). In the absence of the photon injection, we find that spatial non-uniformity has become more pronounced as $L_{y}$ increases. For $L_{y}=20\xi_{0}$, $j_{x}$ is almost parabolic as a function of $y$, but for $L_{y}=40\xi_{0}$, the parabolic nature is somehow deformed at both edges. With photon injection, $j_{x}$ is reduced at the location of the initial hotspot. For $L_{y}=40\xi_{0}$, $j_{x}$ is almost unchanged at the edge. On the other hand, for $L_{y}=20\xi_{0}$, $j_{x}$ outside the hotspot region increases after photon injection and tends to be spatially uniform. The result indicates that the spatial non-uniformity of $j_{x}$ outside the hotspot remains when $L_{y}$ is large.
Figure 11: Time evolution of voltage change $\Delta V$ for $L_{x}\times L_{y}=40\xi_{0}\times 40\xi_{0}$. Various lines present the profiles with different values of $R_{\mathrm{init}}/\xi_{0}$. We take $j_{\mathrm{bias}}/j_{0}=0.28$ and $\kappa=10$.
Based on these observations on the non-uniform current distribution for wider width cases, we examine the time-dependent voltage change $\Delta V$ and the dynamics of the superconducting order parameter for $L_{x}\times L_{y}=40\xi_{0}\times 40\xi_{0}$ to compare it with the previous results with $L_{x}\times L_{y}=40\xi_{0}\times 20\xi_{0}$. The voltage change is shown in Fig. 11. The threshold for successful photon detection is located at $R_{\mathrm{init}}/\xi_{0}=4.6$. The ratio between this threshold and the sample width $40\xi_{0}$ is $0.115$, and is smaller than the ratio $0.145$ in the case of $L_{y}=20\xi_{0}$ (the threshold is located at $R_{\mathrm{init}}/\xi_{0}=2.9$). Thus, the increase in the threshold is weaker than linear, even when we linearly increase the sample width. Near the threshold value, the oscillation of $\Delta V$ is still observed, although the signature of the oscillation is less clear than in the case of narrower width. This oscillation disappears as we increase $R_{\mathrm{init}}$, and this behavior is the same as that in the narrower-width case. Our observation, that the dependence of the threshold on $L_{y}$ is weaker than linear and the vortex–antivortex pair formation is a possible mechanism of photon detection near the threshold, is very helpful for the realization of SWSPD.
Figure 12: Dynamics of the amplitude and phase of the superconducting order parameter for $R_{\mathrm{init}}/\xi_{0}=4.7$ and $L_{x}\times L_{y}=40\xi_{0}\times 40\xi_{0}$. We take $t/t_{0}=312,376,402$ from top to bottom. In the left panels, the contour lines are plotted in the same manner as in Fig. 6(a).
The dynamics of the order parameter is presented in Fig. 12. We find that a vortex–antivortex pair is also formed in the initial hotspot area [see Figs. 12(a)(i) and 12(b)(i)], and then the antivortices penetrate the superconductor from the lower edge [see Figs. 12(a)(ii) and 12(b)(ii)]. After that, vortices and antivortices penetrate from the upper and lower edges repeatedly [see Figs. 12(a)(iii) and 12(b)(iii)]. The motion of the penetrating vortices and antivortices is more significant than that of the vortex–antivortex pair. As shown in Fig. 10, $j_{x}$ near both upper and lower edges of the sample after photon injection is almost independent of the sample width $L_{y}$, even though $j_{x}$ before photon injection is more non-uniform for $L_{y}=40\xi_{0}$ than for $L_{y}=20\xi_{0}$. This result suggests that the significance of the motion of the penetrating vortices and antivortices is not directly related to the non-uniformity of the current density in wider strips.
III.6 Estimation of incident photon energy
[…]
IV Conclusion
We have performed the analysis of $\Delta V$ and $\tilde{\Delta}(x,y)$ on the basis of the TDGL and Maxwell equations to elucidate the photon detection mechanism of SSPD. For $\kappa=10$ and $R_{\mathrm{init}}$ larger than the threshold value, we have successfully observed the voltage change, leading to photon detection. Near the threshold value, we have observed the formation of a vortex–antivortex pair and the subsequent dynamics of the vortex and antivortex. The oscillatory behavior of $\Delta V$ characterizes the penetration of vortices and antivortices from outside of the superconductor. For large-$R_{\mathrm{init}}$ values where $\Delta V$ is smooth as a function of time, the formation mechanism of the normal region is hotspot-like. In the case of $\kappa$ close to the transition point from type-II to type-I superconductor, the oscillatory behavior of the voltage change is also observed, and this is also due to penetrating vortices (antivortices). Our results are qualitatively consistent with experimental results in the sense that a near-critical $R_{\mathrm{init}}$ value provides a reasonable value of $E_{\mathrm{photon}}$. We hope that our analysis serves as a guideline for device fabrication.
Our numerical results and conclusions are somehow different from the previous works [24, 25]. Compared with the results in Ref. [24], even for large-$\kappa$ values, we confirm the usual hotspot mechanism, if $R_{\mathrm{init}}$ is large enough. In Ref. [24] with a small $\kappa$ value, the injection of sufficiently large photon energy generates vortex–antivortex pairs, and this is different from our results. Our model settings and parameter selection are much more systematic. Compared with the results in Ref. [25], the ratio between $R_{\mathrm{init}}$ and $\xi_{0}$ for photon detection due to the presence of vortex–antivortex pairs is different. In Ref. [25], the authors treated the case in which $R_{\mathrm{init}}$ is comparable to $\xi_{0}$, and did not focus on the threshold of $R_{\mathrm{init}}$. On the other hand, there exists the threshold value of $R_{\mathrm{init}}$ in our case, and their comparable setting does not show the finite amount of voltage change. In Ref. [25], the authors focused on the threshold bias current above which the superconducting state becomes unstable, and thus did not pay particular attention to the threshold of $R_{\mathrm{init}}$.
In this paper, we treated the initial process of photon absorption as heating of a fixed region. We did not consider expansion of the initial hotspot region and subsequent heat relaxation of the hotspot into the heat sink, but we obtained a qualitatively reasonable result. We believe that these extra processes do not change our final results.
One of the open questions is about the origin of the response time. We have presented that our response time obtained for $R_{\mathrm{init}}/\xi_{0}=3.0$ is consistent with the experimental results. However, we do not yet understand what determines the magnitude of the response time. This is a future work.
Acknowledgements.
We acknowledge Sanshiro Hatakeyama, Taro Yamashita, Shigehito Miki, Masahiro Yabuno, and Hirotaka Terai for fruitful discussion and providing us with useful information. We acknowledge financial support from KAKENHI projects No. JP23K17323 and No. JP24K17000. HM also acknowledges support from CSIS, Tohoku University.